# Statistical Principles for Platform Trials

### v4

@@ Abstract
While within a clinical study there may be multiple doses and endpoints, across different studies each study will result in either an approval or a lack of approval of the drug compound studied. The term False Approval Rate (FAR) is the term this paper utilizes to represent the proportion of drug compounds that lack efficacy incorrectly approved by regulators. (In the U.S., compounds that have efficacy and are approved are not involved in the FAR consideration, according to our reading of the relevant U.S. Congressional statute).
While Tukey’s (1953) Error Rate Familywise (ERFw) is meant to be applied within a clinical study, Tukey’s (1953) Error Rate per Family (ERpF), defined alongside ERFw, is meant to be applied across studies.
We show that controlling Error Rate Familywise (ERFw) within a clinical study at $5\%$ in turn controls Error Rate per Family (ERpF) across studies at 5-per-100, regardless of whether the studies are correlated or not. Further, we show that ongoing regulatory practice, the additive multiplicity adjustment method of controlling ERpF, is controlling False Approval Rate (FAR) exactly (not conservatively) at 5-per-100 (even for Platform trials).
In contrast, if a regulatory agency chooses to control the False Discovery Rate (FDR) across studies at $5\%$ instead, then this change in policy from ERpF control to FDR control will result in incorrectly approving drug compounds that lack efficacy at a rate higher than 5-per-100, because in essence it gives the industry additional rewards for successfully developing compounds that have efficacy and are approved. Seems to us the discussion of such a change in policy would be at a level higher than merely statistical, needing harmonizsation/harmonization. (In the U.S., policy is set by the Congress.)
@@
1 An motivating example: the HEALEY ALS Platform Trial
Platform trials represent an emerging clinical trial design gaining popularity for their ability to conduct simultaneous testing of different treatments. More significantly, they enable the continuous incorporation of new treatments during the study, facilitating continuous learning. While platform trial design enhances resource efficiency and improves the likelihood of study participants receiving active treatments rather than placebos, the novelty of this design approach leaves ample room for exploration within the field of clinical trials, particularly regarding statistical principles due to the variety of potential outcomes (Koenig et al.,, 2023).
While the excitement surrounding platform trials is undeniable, they do come with several limitations. For instance, platform trials are not universally applicable to all diseases. Meticulous planning is also necessary to ensure their success (Koenig et al.,, 2023).. Nevertheless, despite these challenges, there have been instances of successful implementations of platform trials. One such case is the HEALEY ALS Platform Trial, sponsored by Massachusetts General Hospital.
@@ 1 An motivating example: the HEALEY ALS Platform Trial
ALS is a progressive and fatal neuromuscular disease with the majority of ALS patients dying within 2 to 5 years of receiving a diagnosis. In the HEALEY ALS Platform Trial, the primary outcome is the ALS Functional Rating Scare-Revised (ALSFRS-R) score, which measures 12 aspects of physical functionality. It consists of 12 questions, each targeting a specific aspect of functional capability, with individual scores ranging from 4 (normal function) to 0 (complete loss of function). Thus, the aggregate score can span from a maximum of 48, indicating full functionality, to a minimum of 0, reflecting total loss of physical function. For a repeatedly measured continuous variable such as ALSFRS-R score, Mixed Models Repeated Measures (MMRM) analyses are typically employed.
A significant advantage of a platform trial is the comparison of patients receiving each active treatment to placebo patients shared from other treatment arms (regimens). The sharing of placebo patients (which we will call controls) from other treatment arms (regimens) allows for fewer total participants, and greater allocation of participants to the active treatment within each regimen, relative to traditional designs with 1:1 randomization. However, when employing models like MMRM to analyze repeatedly measured continuous variables such as the ALSFRS-R score, additional considerations emerge beyond the typical variability in estimated efficacy due to patient-to-patient differences within each treatment group. Concerns arise specifically regarding the influence of shared control samples. An overachieving shared control group could inadvertently suggest efficacy across all compounds in a platform trial, presenting challenges for sponsors. Conversely, an underperforming shared control group might lower the bar for all compounds to pass regulatory scrutiny too easily. This article addresses the necessity of planning for a sufficient total sample size to address these concerns while accommodating various randomization ratios that satisfy all stakeholders: patients, regulators, and sponsors.
2 False Approval Rate (FAR) in a regulatory setting
For regulatory agencies like the European Medical Agency (EMA) and the United States Food and Drug Administration (FDA), the purpose of Type I error rate control is to control the rate of approving a treatment (e.g., drug) when it is not efficacious, which we refer to as False Approval Rate (FAR) control. To balance patient benefits and risks, the US FDA requires pharmaceutical companies to provide evidence of a new drugs’ safety and efficacy before it is approved for public use, for which the FDA permits a maximum of five incorrect approvals for every 100 ineffective compounds tested, effectively a 5-per-100 rate. To achieve this, the FDA mandates controlling familywise error rate at $5\%$ within each study. In this paper, we refer to each clinical trail as a “study” and we will show that across studies, FAR is controlled, regardless of whether the studies are independent or correlated in families of platform trials.
2.1 Two layers of families and error rates
@@
In general, when there are decisions to be made both within and across studies, multiple layers of families may exist. For instance, Ding et al., (2018) has a lower layer family comprising three confidence intervals for each single nucleotide polymorphism (SNP), assessing its efficacy across three possible genotypes. At a higher layer, there are confidence statements across all the SNPs in a genome-wide association study (GWAS), evaluating whether each SNP is linked to the phenotype, thus demonstrating the efficacy of a new treatment in comparison to a control. As another example, Han et al., (2022) has a lower layer family of three confidence intervals for each biomarker value, assessing the efficacy of a new treatment versus a control within three distinct groups: the marker-positive $g^{+}$ subgroup, marker-negative $g^{-}$ subgroup, the overall population $\{g^{+},g^{-}\}$. At a higher layer, there are confidence statements across uncountable many biomarker values.
In the Healey ALS Platform Trial, within each regimen, there exists a family of efficacy assessment on primary outcomes (e.g., ALSFRS-R total score disease progression) and secondary outcomes (e.g., respiratory function). Subsequently the regimens in the platform trial form a higher layer of family of studies, each of which either succeeds or fails in claiming efficacy of the compound compared to the placebo.
@@
When there are multiple layers of families of statistical inferences, error rate control can be applied in layers, to control the incorrect decision rate at each layer as well as for the overall decision-making process.
2.2 Tukey’s three Multiple Comparison error rates
It can perhaps be said that Tukey (1953), which has been reprinted as Tukey (1994), started Multiple Comparisons as a discipline in Statistics. The form of statistical inference that John W. Tukey preferred was confidence interval. He defined the following error rates:
@@ 2.2 Tukey’s three Multiple Comparison error rates
$=\frac{\mbox{number of erroneous statements}}{\mbox{number of comparisons}}$
@@
$=\frac{\mbox{number of erroneous statements}}{\mbox{number of families}}$
@@
$=\frac{\mbox{number of erroneous families}}{\mbox{number of families}}$
@@
Note the now common phrasing familywise error rate transposes the error rate and familywise in Tukey’s original terminology. We believe the intent of the per in error rate per family was to have an indication of “how many?” after that. An error rate of 1 per a family of 20, or equivalently an error rate of 5 per a family of 100 (5-per-100 in short), both are $\mathbf{5\%}$ Tukey’s ordering of words is evocative because saying
  • error rate per a family of $k$ studies indicates control of errors is across studies
  • error rate familywise indicates control of errors is within each study
Therefore, instead of the transposed familywise error rate (abbreviated as FWER), we use Tukey’s original ordering error rate familywise (abbreviating it as ERFw). Also contrary to popular perception, it is not true that the only alternative to ERFw is false discovery rate (FDR). While FDR control may be appropriate for medical image reconstruction, it is well known that FDR control is inappropriate for controlling errors within a clinical trial because it can be manipulated by sponsors (see Section 6 of Finner and Roter (2001)). Which multiple comparison error rate to control depends on the specific application as we will discuss below.
2.3 Error Rate Familywise (ERFw) control Guards regulatory False Approval Rate (FAR)
Within a study (a clinical trial), an incorrect statement about a primary endpoint can lead to an incorrect approval of a compound, and an incorrect statement about a secondary endpoint can lead to an incorrect claim of additional efficacy on the drug label. Since any incorrect statement can lead to a (regulatory) incorrect decision, this has led to controlling error rate familywise (ERFw) within each study since the early 2000s.
Different treatment arms within a platform trial may share a common control arm and therefore are not independent. Across the infinitely many studies in independent trials and in platform trials, the following theorem allows us to show if error rate familywise (ERFw) of every study within each platform trial is set at the 5% level, the aforementioned dependence does not affect regulatory false approval rate being no more than 5-per-100 (or 5%), so long as dependence is short-term (within each platform trial only, while studies in different platform trials remain independent).
@@ 2.3 Error Rate Familywise (ERFw) control Guards regulatory False Approval Rate (FAR)
Let $X_{l},X_{2},\ldots,$ have means $\mu_{1},\mu_{2},\ldots$ variances $\sigma^{2}_{1},\sigma^{2}_{2},\ldots$ and covariance $Cov\{X_{i},X_{j}\}$ satisfying
@@
See Theorem E on page 27 of Serfling, (1980). ∎
@@
$X_{i}=\left\{\begin{array}[]{ll}1&\mbox{if a Type I error is made as a result of Study}~{}i,i\in N,\\ 0&\mbox{otherwise}.\end{array}\right.$ (2.3)
@@
Let $m$ = $|j-i|$, then $\rho_{m}=0,~{}m=101,102,\ldots$ for platform trials. Therefore the series $\sum_{1}^{\infty}\rho_{i}$ is convergent. The variance convergence condition is also satisfied since $\sigma^{2}_{i}=\mu_{i}(1-\mu_{i})\leq 1$ and $\sum_{i=1}^{\infty}(\log i)^{2}/i^{2}$ is a convergent series.
@@
This additive multiplicity adjustment (2.4) for controlling ERpF is an exact equality regardless of the dependency structure among the $X_{i}$ variables, whether they are independent, positively dependent, or negatively dependent. It should in no way be confused with the Bonferroni inequality multiplicity adjustment for the purpose of controlling ERFw.
Moreover, Theorem 2.1 assures that as the number of studies ($n$) grows infinitely large, the proportion of studies incurring Type I errors ($V/n$), i.e., ERpF, will converge to the probability of a Type I error in any single study (i.e., ERFw)—$5\%$ in this case. In other words, by the additive multiplicity adjustment (on line 360 of Food and Drug Administration, (2023)), controlling the Type I error rate familywise (ERFw) of each study exactly at 5% leads to controlling the error rate per family (ERpF) at exactly 5% for a family of $n$ studies with in-efficacious compounds (which we phrase as 5-per-100). More generally, regardless of whether studies in the families are independent or dependent, having an error rate per family no more than $5\%\times n$ for a family of $n$ studies means to expect no more than one incorrect approval per a family of 20 studies, five incorrect approvals per a family of 100 studies, etc. Such precision in error control exemplifies the regulatory commitment to balancing risk and benefit for patients, challenging the perception suggested by Bai et al (2020) Bai et al., (2020).
3 Stability of FAR and Conditional Assessment
While reassuring that a 5% ERFw within each study results in an ERpF of 5% across all the studies, there is a concern regarding the stability of ERpF across platform trials. Lines 365 and 366 in Section F, “Multiplicity”, of chapter III , “Considerations On Design and Analysis,” in the Food and Drug Administration (2023) guidelines, specifically state that
@@ 3 Stability of FAR and Conditional Assessment
This section provides a clear demonstration of how to meet this requirement.
Note that it is not whether the unconditional or the conditional error rate should be controlled, as not one error rate control is more meaningful than the other. Both should be assessed because, from elementary statistics, only if one knows the marginal (unconditional) as well as the conditional can one sees the whole picture, i.e. the joint distribution. Thus, think of the platform-specific conditional error rate calculation we describe in this section as within the controlled unconditional error rate over the infinitely many studies over time, conditional on the specific number of studies for a specific platform trial (5 say), with a specific value for the sample mean of the shared control (one SE below its true value say).
3.1 Conditional Stability of FAR is desirable
An error rate represented as an expectation reflects the average value in the probability distribution of errors made. Although an unconditional error rate might be controlled at $5\%$, actual error rates can vary significantly among different families of studies, depending on the characteristics of studies within each family. For example, some families may experience error rates ranging broadly from $1\%$ to $9\%$, while others might have more stable rates, fluctuating only between $4\%$ and $6\%$. Thus, it’s important to consider not only the unconditional (expected) error rate but also the variability of the actual error rates, conditioned on the specific characteristics of the studies within each family.
Let $N$ represent the subset of index $\{1,\ldots,k\}$ for compounds that lack efficacy. For each $i\in N$, let $Z_{i}$ be the indicator that the $i^{th}$ confidence interval erroneously shows efficacy with $P(Z_{i}=1)=\alpha$. Define $V^{*}$ as the sum of these indicators:
$V^{*}=\sum_{i\in N}Z_{i}.$
The variance of $V^{*}$, given by
$Var(V^{*})=\sum_{i}Var(Z_{i})+\sum_{i\neq j}Cov(Z_{i},Z_{j})$
where $Var(Z_{i})=\alpha(1-\alpha)$, can be calculated once the covariances $Cov(Z_{i},Z_{j})$ for $i\neq j$ are known. This section will discuss two specific concerns on variability of $V^{*}$ associated with sharing a control in a platform trial.

### v5

# Principles of Conditionality and Layering of Error Rates with Application to Platform Trials

@@ Abstract
There has been a misconception that only one type of error rate control is necessary in clinical trials, leading to debates over whether to prioritize Familywise Error Rate (FWER) or False Discovery Rate (FDR). This misconception has led to misleading statements about FWER control and proposals to shift towards FDR control, which could be manipulated by the industry.
In reality, since the early 2000s, biopharmaceutical statistics have implicitly applied two layers of Type I error rate control. This aligns with Tukey’s 1953 invention of Error Rate per Family (ERpF) for controlling error across studies, while FWER applies within each study. Our paper clarifies this layering, using Platform trials to demonstrate the verifiable conditions needed across studies for the FDA to fulfill its regulatory mission.
We show that controlling FWER within a study at $5\%$ inherently controls ERpF across studies at 5-per-100, regardless of study correlations. This supports current regulatory practices that protect public health while fostering innovation. We also address concerns about ERpF stability in Platform trials, where shared controls introduce dependencies. By applying the Conditionality Principle and utilizing an innovative Shiny app, we explore how correlations impact ERpF variability, providing deeper insights for informed decision-making.
Our findings, supported by principles like Layering of Error Rate Controls and the No-gaming Principle, are particularly relevant as Platform trials gain popularity for their efficiency in testing multiple treatments simultaneously.
Correspondence: Xinping Cui (xinping.cui@ucr.edu)
@@
1 Introduction
Clinical trials are, and have long been, an integral part of biomedical and pharmaceutical research. But as clinical study designs continue to evolve, the complexity of these trials has increased dramatically. Modern frameworks, including adaptive designs, platform trials, and multistage testing, introduce additional layers of structure, dependence, and decisions. Many classical methods of clinical trial analysis were developed under simplifying assumptions that are no longer appropriate in these settings. When applied naively, such classical procedures can produce inferences that are logically inconsistent and may fail to maintain error rate control.
@@ 1 Introduction
Principle 1.1 (The Conditionality Principle).
See Berger (1985) page 31 and Casella and Berger (1990) page 268 for details.
Separate from these inferential concerns, there is the regulatory goal of controlling False Approval Rate (FAR). To make explicit the multi-level nature of regulatory decision making, we propose the Principle of Layering of error rates, which emphasizes that error rates should be controlled at each layer of decision making which is detailed in Section 2.
[…]
2 Layering of Error Rates
In this section, we describe the layered error-rate framework underlying current regulatory practice. While elements of this framework have appeared in prior work, we present it here in a unified form that distinguishes error control within studies from error control across studies.
2.1 Incorrect Approval Rate (IAR) in a regulatory setting
For regulatory agencies like the European Medical Agency (EMA) and the United States Food and Drug Administration (FDA), the primary purpose of Type I error rate control is to limit the rate of approving a treatment lacking efficacy (e.g., drug), which we refer to as Incorrect Approval Rate (IAR) control. In the United States, the FDA currently permits at most five incorrect approvals per one hundred ineffective compounds studied, corresponding to an IAR of 5-per-100. Rather than attempting to control this error rate directly across various study designs, the FDA instead mandates that each individual study controls its familywise error rate at $5\%$ within each study. In enforcing this mandate, control of IAR at $5\%$ across studies is also achieved. To understand what this means, it is important to define the different layers of error-rate control.
2.1.1 Two layers of families and error rates
In many modern inferential settings, when there are decisions to be made both within and across studies, two layers of families may exist. As such, there are two different error rates we need to consider:
@@ 2.1.1 Two layers of families and error rates
These layers correspond to different stages at which information is revealed and decisions are made. For instance, Ding et al., (2018) has a group of inferences (lower layer family) comprising four simultaneous confidence intervals for each single nucleotide polymorphism (SNP), assessing treatment efficacy across three possible genotypes. At the higher layer, there are a panel of inferences comprising confidence statements across all SNPs on a genome-wide scale, evaluating whether each SNP is “predictive” of treatment efficacy. As another example, Han et al., (2022) has a lower layer family of three confidence intervals for each biomarker value, assessing the efficacy of a new treatment versus a control within three distinct groups: the marker-positive $g^{+}$ subgroup, marker-negative $g^{-}$ subgroup, the overall population $\{g^{+},g^{-}\}$. At the higher layer, there are confidence statements across uncountable many biomarker values.
In platform trials, within each regimen, there exists a family of efficacy assessment on primary and secondary outcomes. For example, the HEALEY ALS Platform Trial has the primary outcomes (e.g., ALSFRS-R total score disease progression) and secondary outcomes (e.g., respiratory function). Subsequently the regimens in the platform trial form a higher layer of family of studies, each of which either succeeds or fails in claiming efficacy of the compound compared to the placebo.
@@
When there are multiple layers of families of statistical inferences, error-rate control can be applied in layers, to control the incorrect decision rate at each layer as well as for the overall decision-making process.
Because decision-making occurs sequentially at multiple inferential levels—both within and across studies — separate error-rate controls are needed at each layer, necessitating layering of error rates.
2.1.2 Tukey’s three Multiple Comparison error rates
Tukey (1953; reprinted in 1994) formalized multiple comparisons by framing inference in terms of confidence intervals and introducing several distinct error rates.
@@ 2.1.2 Tukey’s three Multiple Comparison error rates
$\frac{\mbox{number of erroneous statements}}{\mbox{number of comparisons}}$
@@
$\frac{\mbox{number of erroneous statements}}{\mbox{number of families}}$
@@
$\frac{\mbox{number of erroneous families}}{\mbox{number of families}}$
@@
2.1.3 Error Rate Familywise (ERFw) control guards regulatory Incorrect Approval Rate (IAR)
Within a study (a clinical trial), an incorrect statement about a primary endpoint can lead to an incorrect approval of a compound, and an incorrect statement about a secondary endpoint can lead to an incorrect claim of additional efficacy on the drug label. Since any incorrect statement within a study can lead to a (regulatory) incorrect decision, this has led to controlling error rate familywise (ERFw) within each study since the early 2000s.
Within a platform trial, different treatment arms may share a common control arm and therefore are not independent. This dependence is short-term, meaning that studies are correlated within the groups of studies that share data or analysis infrastructure, while studies without such sharing are independent. Across the infinitely many studies in independent trials and in platform trials, the following theorem allows us to show if the ERFw of each study is set at 5% level, the aforementioned dependence does not inflate the regulatory IAR beyond 5-per-100 (or 5%), as long as dependence is short-term (within each platform trial only, while studies in different platform trials remain independent).
We now establish an unconditional guarantee for error-rate control across studies, demonstrating that layered error-rate control remains valid in the long run even under short-term dependence.
@@ 2.1.3 Error Rate Familywise (ERFw) control guards regulatory Incorrect Approval Rate (IAR)
Let $X_{1},X_{2},\ldots,$ have means $\mu_{1},\mu_{2},\ldots$ variances $\sigma^{2}_{1},\sigma^{2}_{2},\ldots$ and covariance $Cov\{X_{i},X_{j}\}$ satisfying
@@
See Theorem E on page 27 of Serfling(1980). ∎
@@
[…]
@@
Let $m$ = |ji||j-i|, then $\rho_{m}=0,~m=101,102,\ldots$ for platform trials. Therefore the series $\sum_{1}^{\infty}\rho_{i}$ is convergent. The variance convergence condition is also satisfied since $\sigma^{2}_{i}=\mu_{i}(1-\mu_{i})\leq 1$ and $\sum_{i=1}^{\infty}(\log i)^{2}/i^{2}$ is a convergent series.
@@
This additive-$\alpha$ multiplicity adjustment (2.4) for controlling error rate of 5 per a family of n studies is an exact equality regardless of the dependency structure among the $X_{i}$ variables, whether they are independent, positively dependent, or negatively dependent. It should in no way be confused with the Bonferroni inequality multiplicity adjustment for the purpose of controlling ERFw.
[…]
We include this result to clarify while long-run regulatory guarantees remain valid, they do not by themselves ensure coherent inference when decisions are interpreted conditionally on the realized experiment.
Principle 2.2 (The No-gaming Principle).
A regulatory approval system should not allow sponsors to game it.
Principle 2.2 is a fundamental policy requirement for any regulatory approval system. For example, FDR control is not used to control ERFw within a clinical trial because (as already mentioned) it can be manipulated by sponsors. Similarly, if the requirement for ERpF control across studies were merely to control $E[V]$ at $5\%\times n$, then an individual sponsor attempting to gain approval for its compound ${Rx}^{1}$ can game the system by conducting a Platform trial of its own (artificially) adding compounds ${Rx}^{i},i=2,3,4$ which are of no commercial value, controlling the ERFw of ${Rx}^{1}$ at 17% while controlling the ERFw of each of ${Rx}^{i},i=2,3,4$ at 1%, gaming the system. To adhere to the No-gaming Principle, industry should continue with the current practice of the additive-$\alpha$ adjustment, controlling the ERFw of each study at $\alpha$.
3 Applying the Conditionality Principle in Practice
We have previously established that strong ERFw control within each study ensures valid unconditional control of the IAR across studies over time. However, in platform trials, regulators and stakeholders may also be concerned with behavior of the IAR within a given platform. This concern is reflected in FDA (2023) guidance. In section F (“Multiplicity”) of chapter III (“Considerations On Design and Analysis,”) of Food and Drug Administration (2023) guidance document, it is explicitly stated that
@@ 3 Applying the Conditionality Principle in Practice
It emphasizes that regulators are concerned not only with the expected number of false approvals (i.e. the unconditional IAR), but also with the variability and distribution of IAR within a given design (i.e., conditional IAR). A complete understanding of error behavior requires consideration of both the marginal (unconditional) and conditional perspectives, as together they characterize the joint distribution of IAR.
Motivated by this perspective, we examine platform-specific conditional IAR, calculated under realized design features such as the number of regimens and the observed shared control outcome, while preserving unconditional IAR control across an infinite sequence of studies.
3.1 Platform Trials
Platform trials represent an emerging clinical trial design that has gained increasing popularity due to their ability to evaluate multiple treatments simultaneously within a unified framework. This design enhances resource efficiency and increases the likelihood that participants receive active treatments rather than placebos. However, the relative novelty and structural complexity of platform trials leave substantial room for further methodological development, particularly with respect to foundational statistical principles and error rate control in the presence of multiple, interdependent comparisons(Koenig et al.,, 2023).
Because statistical inference and regulatory decisions are made after the studies are conducted, platform trials provide a concrete setting in which the conditionality principle is not merely philosophical but operationally essential. In such trials, inference and regulatory decision-making proceed conditional on multiple layers of realized information, including the number of treatment arms, enrollment patterns, interim adaptations, and the performance of the shared control group.
A prominent example is the HEALEY ALS Platform Trial, a double-blind, placebo-controlled study evaluating multiple investigational therapies for amyotrophic lateral sclerosis (ALS). In this trial, participants are randomized across treatment regimens, each compared against a shared control group. The shared control structure induces statistical dependence across treatment comparisons and naturally give rise to conditioning event that are central to inference and error rate assessment. These features motivate a formal conditional assessment of IAR control within a platform trial, which we develop in the next section.
3.2 Conditional assessment of IAR control
Applying Principle 1.1, inference within a platform trial should be interpreted conditional on realized experiment, including the observed shared outcome and the dependence structure induced by control sharing. In practice, the relevant condition events include the number of regimens, the observed shared control mean, and the correlation structure across treatment comparisons.
To motivate this conditional perspective, consider Fieller’s confidence interval for the ratio of two Normal means(Fieller,, 1954). Unconditionally, the interval achieves exact $(1-\alpha)$ coverage across infinitely many replications. However, the form of Fieller’s confidence interval depends on the realized data; it may be finite, infinite (the entire real line), the complement of an interval, or empty (see Casella and Berger (1990), page 459). Conditional on the interval being the entire real line, it conveys no information about the parameter; conditional on it being empty, the effective confidence is 0%. These examples illustrate that while unconditional guarantees remain valid, they do not fully describe the evidential strength in the realized experiment. A complete assessment therefore requires consideration of both the marginal (unconditional) properties and the conditional behavior given the observed data.
Platform trials exhibit this same phenomenon. While a 5% ERFw within each study guarantees an unconditional IAR of 5% across an infinite sequence of studies, the evidential strength and error behavior within a particular platform depend on the realized shared control outcome and study configuration. Accordingly, we recommend assessing the conditional IAR within each platform trial, given the observed shared control outcome and realized design features, to ensure that it does not deviate materially from its unconditional guarantee.
3.3 Conditional Stability of IAR is desirable
We now study the distribution of erroneous efficacy claims within a single platform trial under the shared control structure. Although unconditional IAR control constrains the expected error rate across repeated experimentation, conditional error behavior may vary across realized platform configurations. Our goal is to characterize this variability formally.
Consider $k$ studies in a platform trial. Without loss of generality, we define efficacy as $\theta_{i}>0$ and consider one-sided inference. Let
$L=\left\{i\in\{1,\dots,k\}:\theta_{i}=\mu_{i}-\mu_{SC}\leq 0\right\}$
denote the set of ineffective compounds and let $|L|=\ell$. $\mu_{i}$ and $\mu_{SC}$ are true mean of the new compound $i$ and the shared control, respectively. Define
[…]
so that for each $i\in L$, $Z_{i}=1$ corresponds to an erroneous efficacy conclusion (an incorrect approval), The total number of incorrect approvals is therefore
$V^{k}=\sum_{i\in L}Z_{i}.$
The distribution of $V^{k}$ depends on the unknown configuration $L$.
Since neither the identities nor the number of $\theta_{i}\leq 0$ are known, the exact distribution of $V^{k}$ cannot be determined without additional assumptions. In what follows, we will obtain a bound $V^{*}$ that does not depend on $L$, and discuss two specific concerns on variability of $V^{*}$ associated with sharing a control in a platform trial and the need for conditioning in accordance with the Conditionality Principle.